A car with a mass of 500 kg contains 2600 J of kinetic energy. Calculate the speed of the car.



1
Expert's answer
2021-08-29T16:57:30-0400
"KE=\dfrac{1}{2}mv^2,""v=\sqrt{\dfrac{2KE}{m}},""v=\sqrt{\dfrac{2\cdot2600\ J}{500\ kg}}=3.22\ \dfrac{m}{s}."